# Magic tricycles: efficient magic state generation with finite block-length quantum LDPC codes

### v1

@@ Appendix B Cup products and transversal CCZ action
The depth $18$ circuits are triple-sector circuits constructed by choosing any preorientations with $|\alpha_{I}^{in}|=1$, $|\alpha_{I}^{out}|=3$, $|\alpha_{II}^{in}|=1$, $|\alpha_{II}^{out}|=3$, $|\alpha_{III}^{in}|=1$, $|\alpha_{III}^{out}|=3$. Then, for fixed $p^{i}$, $|p^{-1}\alpha_{i}^{in}\cap q^{-1}\alpha_{j}^{out}|=1$ for exactly $3$ different $q^{j}$. For each such $q^{j}$, $|q^{-1}\alpha_{j}^{in}\cap r^{-1}\alpha_{k}^{out}|=1$ in turn for exactly $3$ different $r^{k}$. Thus, each distinct sector-triple $i,j,k$ contributes 99 CCZCCZ gates acting on each qubit in the sectors. Each sector of each code-block is involved in two such triples – for example, sector $I$ of block $1$ is involved in CCZ gates acting as IIIIIII-II-III and IIIIIII-III-II between code-blocks. Each such triple contributes $9$ gates, leading to a total depth $18$ circuit with the structure illustrated in Fig.˜1 c). ∎
Consequently, all tricycle codes in footnote˜2 have depth-$18$ CCZCCZ circuits, constructed from any partition of the form described in the preceding proof. Using similar reasoning, one can show that larger circuit depths are possible. For example, a SS circuit with $|\alpha_{I}^{in}|=4$, $|\alpha_{II}^{out}|=|\alpha_{II}^{in}|=2$, $|\alpha_{III}^{out}|=4$ with no further structure between the preorientations will generically lead to a depth $4\times 2\times 2\times 4=64$ circuit. However, we are of course interested primarily in shorter depth $CCZ$ circuits. By imposing additional relations between the $\alpha$ group-algebra elements, this can be accomplished. We illustrate how this can be done to achieve a depth-4 SS circuit in the proof of the following theorem. Circuits with other attainable depths $8,12,16$ for both SS and TS circuits are constructed using a very similar method.
Theorem B.3.
There exist tricycle codes constructed from weight-$4$ group algebra elements $\alpha_{I},\alpha_{II},\alpha_{III}$ which have depth-$4$ single-sector code-space preserving $CCZ$ circuits.
Proof.
We derive conditions that $\alpha_{I},\alpha_{II},\alpha_{III}$ must satisfy to be compatible with such a circuit. We will consider preorientiations with the single-sector circuit pattern $|\alpha_{I}^{in}|=4$, $|\alpha_{II}^{out}|=|\alpha_{II}^{in}|=2$, $|\alpha_{III}^{out}|=4$ for the $f_{CCZ}$ function in Eq.˜26. Let us explicitly write out the group algebra elements as
$\alpha_{I}^{in}=\alpha_{I}$ $=a+b+c+d$ (27)
$\alpha^{out}_{II}$ $=e+f$ (28)
$\alpha^{in}_{II}$ $=g+h$ (29)
$\alpha^{out}_{III}=\alpha_{III}$ $=i+j+k+l$ (30)
for $a\cdots l\in G$, and $a\neq b\neq c\neq d\$, $e\neq f\neq g\neq h\$, $i\neq j\neq k\neq l$. As discussed in the preceding paragraph, without any additional structure, these partitions would produce a depth $64$ circuit. To be more explicit, this is because for every one of the $4$ group elements in $\alpha_{I}^{in}$ and every one of the $2$ group elements in $\alpha^{out}_{II}$, the group structure implies that there are $p,q\in G$ that make the first intersection term in the expression for $f_{CCZ}$ in Eq.˜26 equal to $1$. This gives a factor of 88 $(p,q)$ pairs from the first term in the expression for $f_{CCZ}$, and the second term similarly gives a factor of 88 $(q,r)$ pairs from considering the intersections of $\alpha^{out}_{III}$ and $\alpha^{in}_{II}$. Therefore, to reduce the circuit depths, we must reduce the number of pairs $p,q$ that make the first intersection term in Eq.˜26 non-zero, and independently reduce the number of pairs $q,r$ that do so for the second intersection term.
Consider imposing the following group relations between the components of $\alpha^{in}_{I}$ and $\alpha^{out}_{II}$:
$af$ $=be$ (31)
$bf$ $=ce$ (32)
$cf$ $=de$ (33)
Then, for any $p,q\in G$ such that $p^{-1}a=q^{-1}e$, the first equation implies that $p^{-1}b=q^{-1}f$ must also be true. Thus, the intersection is made even (i.e, 0mod(2)0\ mod(2)) whenever $a,e$ are involved in a non-trivial intersection of the first term in Eq.˜26 by forcing $b,f$ to simultaneously be involved in the intersection. This reduces the contribution from the first term in Eq.˜26 towards the number of $CCZ$ gates by $2$. Thus, imposing only the very first relation results in a circuit of depth $6\times 8=48$. Similarly imposing the second and third equations each reduce the first factor by $2$ each by pairing group terms involved in non-trivial intersection. Thus, imposing all three relations in Eq.˜31 leads to a circuit of depth $2\times 8$ = $32$. One may be tempted to further impose a fourth relation such as $df=ce$, but a quick calculation shows that satisfying all $4$ equations is only possible if one of the partitions is identically 0. Similarly, one can impose the following relations between $\alpha^{out}_{II}$ and $\alpha^{in}_{III}$:
$ih$ $=gj$ (34)
$jh$ $=gk$ (35)
$kh$ $=gl$ (36)
Each equation reduces the contribution from the second term in Eq.˜26 from $8$ by $2$. Thus, imposing all $6$ equations in Eq.˜31 and Eq.˜34 leads to a final circuit depth of $2\times 2=4$. Finally, existence of tricycle codes with good parameters that are constructed from group-algebra elements satisfying these relations is demonstrated numerically, by sampling group algebra elements that constructively satisfy the above relations. See examples in Table˜2.
The proof of Theorem˜B.3 demonstrates how we find tricycle codes with short-depth circuits. We frame the equations in Eq.˜31 and Eq.˜34 as linear systems over the integer powers of group-generators corresponding to each group element. Then, we row-reduce the system to determine which of the variables $a\cdots l$ are free. These free variables are then randomly sampled from a given group, determining the other involved group elements. We filter sampled codes by rate and distance, using the methods in Appendix˜D. Using a similar ‘pairing’ idea between group elements, one can similarly construct SS circuits and TS circuits of depth $8,12,16$. For example, a SS circuit with depth $8,12,$ or $16$ can be constructed by enforcing a subset of the equations in Eq.˜31 and Eq.˜34. We do not present all the details here, but one enforces a desired number of pairwise relations between involved group elements as in Eq.˜31 and Eq.˜34 to reduce the contribution of terms in Eq.˜26 appropriately. In the LABEL:table::short_code_polys, we provide the partitions used to construct the codes in Table˜2 that have $CCZ$ circuits with depth shorter than $18$.
Table 5: Polynomials of partitions used to construct codes in Table˜2 (listed in the same order) with short-depth CCZCCZ circuit. Polynomials represent elements of the group algebra of $G=\mathbb{Z}_{l}\times\mathbb{Z}_{m}\times\mathbb{Z}_{n}$. A dash indicates that a partition is empty. Codes are constructed by imposing group element equations similar to Eq.˜31 and Eq.˜34 and randomly sets of sampling group algebra elements that satisfy these relations.
$\mathbf{[[N,K,(D_{X},D_{Z})]]}$ $\mathbf{(l,m,n)}$ Depth $\mathbf{a_{in}}$ $\mathbf{a_{out}}$ $\mathbf{b_{in}}$ $\mathbf{b_{out}}$ $\mathbf{c_{in}}$ $\mathbf{c_{out}}$
$135,3,\leq 22,8]]$ $(3,3,5)$ 4 $1+yz+y^{2}z^{2}+z^{3}$ _\_ $1+xz^{3}$ $z^{3}+yz^{4}$ _\_ $xy+yz+yz^{2}+x^{2}yz^{3}$
$405,9,(\leq 27,10)]]$ $(3,5,9)$ 44 $1+yz^{2}+z^{4}+z^{8}$ _\_ $y^{4}z+z^{6}$ $y^{4}+y^{4}z^{8}$ _\_ $y^{3}+yz+y^{3}z^{2}+y^{4}z^{5}$
$135,3,(\leq 27,9)]]$ $(3,3,5)$ 88 $1+xz^{3}+x^{2}z^{4}+z^{4}$ _\_ $yz+y^{2}z^{3}$ $x^{2}yz^{4} _\_ y+y$$yz+y^{2}z^{2}+z^{4}$
$405,9,(\leq 29,11)]]$ $(3,5,9)$ 88 $1+yz^{7}+y^{2}z^{3}+y^{4}z^{6}$ _\_ $yz^{2}+z^{7}$ $y^{2}z^{8}+xy^{4}z^{2} _\_ x+x$$xyz^{4}+xy^{3}z+xy^{4}z^{5}$
$405,9,(\leq 37,14)]]$ $(3,5,9)$ $12$ $1+z+x^{2}y^{2}z^{5}+x^{2}y^{3}z^{6}$ _\_ $xz^{8}+xy^{4}z^{2}$ $x^{2}y^{3}z+x^{2}y^{3}z^{2}$ _\_ $z^{5}+yz^{2}+y^{2}z^{8}+y^{4}z^{8}$
$405,9,(\leq 27,11)]]$ $(3,5,9)$ $12$ $1+z3$ $x^{2}y^{3}z^{3}+x^{2}y^{4}z^{5}$ $z^{2}+y^{3}z^{7}+y^{4}$ $y^{4}z^{8}$ $x^{2}y+x^{2}y^{2}z^{2}+x^{2}y^{3}z^{4}$ $x^{2}yz^{4}$
$375,15,(\leq 15,10)]]$ $(5,5,5)$ $16$ $1+y^{2}$ $x^{3}yz^{4}+x^{4}y^{4}$ $x^{2}z^{2}+x^{3}y^{3}z^{3}$ $x^{3}y^{2}z^{2}+x^{3}y^{4}z^{2}$ $y^{2}z^{3}+y^{4}z^{3}$ $xy+x^{2}y^{4}z$
Appendix C Logical circuit optimization
The logical connectivity is determined by the cohomology invariant $f_{CCZ}$ function restricted to a basis of logical $X$-type operators for the three code-blocks. In cohomological terms, this is a basis for the first cohomology group $H^{1}(C)$ of the associated co-chain complex. Denote a choice of these bases as $\{[l^{1}_{i}]\}_{i=1\cdots K},\{[l^{2}_{j}]\}_{j=1\cdots K},\{[l^{3}_{k}]\}_{k=1\cdots K}$ where the $l_{i}$ are representative elements of $\mathrm{ker}(H_{Z})$ and $\cdot$] denotes an equivalence class modulo elements of $\mathrm{im}(H_{Z}^{T})$. We then construct a $K\times K\times K$ binary $3$-tensor corresponding to the restriction of $f_{CCZ}$ to the logical operators:
$T^{log}_{ijk}=f_{CCZ}(l^{1}_{i},l^{2}_{j},l^{3}_{k})\in\mathbb{F}_{2}$ (37)
$(M^{1}\otimes M^{2}\otimes M^{3})\cdot T^{log}=\mathbb{I}^{r\times r\times r}\ \mathrm{mod}(2)$ (38)
[…]
Unfortunately, finding the subrank $r$ and the associated bases $M^{1},M^{2},M^{3}$ is a problem that is widely believed to be NP-Hard, even over the real numbers [48]. In this work, we find sub-optimal solutions using a Mixed-Integer-Programming (MIP) method – we use the Gurobi Python package for this task [38]. We first find a set of initial logical operators numerically using row-reduction over $\mathbb{F}_{2}$. For fixed $r$ we then treat the $3rK$ entries of $M^{1},M^{2},M^{3}$ as binary variables for the integer program, and enforce the condition in Eq.˜39 as a set of integer constraints which involve another $2r^{3}$ many slack variables. The complexity of the integer program therefore grows dramatically with $r$. We iterate from $r=0$ to increasing values of $r$ until the MIP solver converges or is unable to find a feasible solution within time constraints. This leads to the reported values of $K_{CCZ}$ in Sec.˜2.2. An important direction for future work is to find tailored optimization heuristics for the binary tensor subrank problem to extract larger values of $K_{CCZ}$, signifcantly improving the throughput of $\overline{CCZ}$ type magic states from the hypergraph magic states produced by our circuits.
Appendix D Determining code distances
@@
The first method is a Mixed-Integer-Program (MIP). This involves first using Gaussian row-reduction over $\mathbb{F}_{2}$ to find a basis of logical $X$ and $Z$ operators $L_{X}$ and $L_{Z}$ independently for a given code. One then formulates the low-weight logical operator search as a mixed-integer-program by iterating over the basis of logical operators. For example, to find $d_{X}$, for the $i^{th}$ logical operator $l_{i}\in L_{Z}$, we minimize the hamming weight of candidate bitstring solutions $b$ under the constraints $H_{Z}b=0\ mod(2)$ and $b^{T}l_{i}=1\ mod(2)$ where the latter constraint enforces that $b$ corresponds to an $X$-type logical operator that anticommutes with the $i^{th}$ logical $Z$ operator. The minimum weight logical operator is then found after iterating over all $l_{i}$ if the solver converges within time constraints. We used the Gurobi [38] package in Python for this task.

### v2

@@ Appendix B Single-shot guarantees
First, note that it is easy to extend Theorem˜B.4 to when the classical seed codes are different, though this generally results in the bases having different soundness parameters. Then, recognizing that a tricycle code is just a 4D $R$-homological product code in which one of the seed codes is a degree-0 chain complex [47], one can check that the unaffected basis remains sound with the claimed $t=\min_{i}d_{i}$ and the claimed $g$. As mentioned in Sec.˜2.3, one can modify a minimum-weight decoder to decode the metacheck matrix $H_{\text{meta}}$ in such a way that $d_{ss}=\infty$. Then, combining Theorem 1 of [22], Theorem˜A.2, and Theorem˜B.4, we obtain the corollary. ∎
As with Corollary˜A.3.1, this result can be readily extended to disconnected codes. Even though Corollary˜B.4.1 only applies to codes where $|N|$ is odd, we expect, based on our numerics in Sec.˜2.3 and Sec.˜2.4 that codes with even $|N|$ are sound. In fact, many of our codes in LABEL:table::code_polys are connected with even-order intersection subgroups.
Appendix C Determining code distances
@@
The first method is a Mixed-Integer-Program (MIP). This involves first using Gaussian row-reduction over $\mathbb{F}_{2}$ to find a basis of logical $X$ and $Z$ operators $L_{X}$ and $L_{Z}$ independently for a given code. One then formulates the low-weight logical operator search as a mixed-integer-program by iterating over the basis of logical operators. For example, to find $d_{X}$, for the $i^{th}$ logical operator $l_{i}\in L_{Z}$, we minimize the hamming weight of candidate bitstring solutions $b$ under the constraints $H_{Z}b=0\ \mathrm{mod}(2)$ and $b^{T}l_{i}=1\ \mathrm{mod}(2)$ where the latter constraint enforces that $b$ corresponds to an $X$-type logical operator that anticommutes with the $i^{th}$ logical $Z$ operator. The minimum weight logical operator is then found after iterating over all $l_{i}$ if the solver converges within time constraints. We used the Gurobi [41] package in Python for this task.